OpenStudy (anonymous):

the temperature was 65 degrees at daybreak. then it dropped 2 degrees per hour until dusk. if x is the number of hours since daybreak, this equation represents the temperature during the day: y=-2x+65 OR y=65- 2x. use the equation to predict the temperature 6 hours later

OpenStudy (anonymous):

both eq are same y = -2x+ 65 = 65-2x Put x = 6